# SARChat-Bench-2M: A Multi-Task Vision-Language Benchmark for SAR Image Interpretation

### v1

@@ Abstract
In the field of synthetic aperture radar (SAR) remote sensing image interpretation, although Vision language models (VLMs) have made remarkable progress in natural language processing and image understanding, their applications remain limited in professional domains due to insufficient domain expertise. This paper innovatively proposes the first large-scale multimodal dialogue dataset for SAR images, named SARChat-2M, which contains approximately 2 million high-quality image-text pairs, encompasses diverse scenarios with detailed target annotations. This dataset not only supports several key tasks such as visual understanding and object detection tasks, but also has unique innovative aspects: this study develop a visual-language dataset and benchmark for the SAR domain, enabling and evaluating VLMs’ capabilities in SAR image interpretation, which provides a paradigmatic framework for constructing multimodal datasets across various remote sensing vertical domains. Through experiments on 16 mainstream VLMs, the effectiveness of the dataset has been fully verified, and the first multi-task dialogue benchmark in the SAR field has been successfully established. The project will be released at https://github.com/JimmyMa99/SARChat, aiming to promote the in-depth development and wide application of SAR visual language models.
@@
footnotetext: These authors contributed equally to this work.**footnotetext: Corresponding author. Xiayang Xiao and Haipeng Wang
@@ 1 Introduction
In recent years, deep neural networks, particularly CNNs (Dosovitskiy, 2020) and ViTs (LeCun et al., 1998), have achieved remarkable progress in remote sensing data analysis, significantly enhancing the processing efficiency and analytical accuracy of large-scale datasets. However, existing research primarily focuses on visual feature extraction, while lacking in deep semantic parsing and general reasoning capabilities (Li et al., 2024), limiting model applicability and interpretability in complex scenarios.
With the advancement of Large-Language Models (LLMs) in natural language processing, Vision-Language Models, through integrating generative pre-training and instruction tuning strategies, have demonstrated robust zero-shot learning and generalization capabilities in multimodal interactive tasks Dai et al. (2023). This has inspired researchers to explore the deep integration of visual models with LLMs.
Although preliminary achievements have been made with models like RSGPT (Hu et al., 2023) and GeoChat (Kuckreja et al., 2024), these approaches are primarily designed for natural images and struggle to perform well in SAR applications. SAR images inherently pose significant interpretation challenges due to their scattering imaging mechanisms, characterized by blurred target edges, dispersed speckles, and orientation sensitivity. Meanwhile, existing SAR datasets primarily focus on visual recognition tasks (Kuckreja et al., 2024; Cheng et al., 2022; Zhang et al., 2023), leaving a critical shortage of large-scale, high-quality image-text alignment datasets. Both these intrinsic characteristics and data limitations impede the advancement of VLMs in the SAR domain.
@@
As shown in Figure 1, we present SARChat-2M, a large-scale multimodal conversational dataset for SAR images, and establish SARChat-Bench, a comprehensive multimodal task-oriented benchmark for the SAR domain. The SARChat-2M dataset contains approximately 2 million high-quality SAR image-text pairs across maritime, terrestrial, and urban scenarios, featuring fine-grained semantic descriptions and multi-scale resolutions (0.3-10 meters). Through cross-modal representation learning, the dataset enables multi-task learning capabilities, including image captioning, VQA(Visual Question Answering), visual localization, and object detection. The SARChat-2M systematically evaluates vision-language models’ performance through six core tasks: classification, description, counting, localization, recognition, and refering.
The primary contributions of this paper are as follows.
(1) The construction of SARChat-2M, the largest SAR remote sensing instruction-following dataset to date, comprising over 2 million high-quality image-text pairs across multi-scenario task-oriented dialogues, addressing the knowledge scarcity of VLMs in the SAR domain.
(2) The development of SARChat-Bench, a comprehensive SAR domain multimodal benchmark encompassing six core tasks (classification, description, counting, localization, recognition, and refering), enabling systematic evaluation of vision-language models through multi-dimensional assessment metrics.
(3) It pioneers a research paradigm applicable to the SAR field, providing reference ideas for the construction of models in other remote-sensing vertical domains. The methods and processes adopted in data collection, annotation, as well as model training and evaluation in this study have good generality and extensibility.
@@ 2.1 VLMs for Remote Sensing
VLMs are capable of converting images into natural language descriptions and parsing the relationships between objects, demonstrating remarkable performance in tasks such as text-image retrieval, image captioning, and visual question answering. Recently, models like RemoteClip (Liu et al., 2024) have been applied to the field of remote sensing images, primarily focusing on cross-modal retrieval and zero-shot classification. However, these models have not addressed tasks such as image description generation and visual grounding. The RSGPT model has achieved text description and visual question answering for remote sensing images, but it has not expanded to tasks such as classification and detection. The GeoChat model has advanced multi-task conversational processing of high-resolution remote sensing imagery, including scene classification, visual question answering, multi-turn dialogue, visual grounding, and reference object detection. However, these models, including GeoChat, are primarily trained on optical remote sensing data, which limits their application scope. EarthGPT (Zhang et al., 2024) has extended the application of multimodal large language models to the remote sensing field through instruction tuning, but its performance in SAR image multi-task processing still needs improvement. Compared with natural images, the interpretation of SAR images is more challenging, which poses higher demands on the model’s processing capabilities and adaptability.
@@ 2.2 Remote Sensing Vision-Language Datasets
Remote sensing datasets are essential for intelligent interpretation models. Existing datasets such as UCM Captions, Sydney Captions Qu et al. (2016), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), and RSVG (Zhan et al., 2023) provide preliminary resources for studying the correlation between remote sensing images and text, but they are limited in scale and only cover optical images. Although large-scale datasets like MillionAID (Long et al., 2021), FMoW (Christie et al., 2018), and BigEarthNet (Sumbul et al., 2019) exist, they lack text-image pairs. The RS5M dataset (Zhang et al., 2023), containing 5 million image-text pairs, is still limited to optical images. The MMRS-1M dataset (Zhang et al., 2024), which covers optical, infrared, and SAR modes, has a very low proportion of SAR image-text data. Therefore, this paper constructs the SARChat-2M dataset, which focuses on SAR images and contains over 2 million image-text pairs, covering tasks such as classification, detection, caption generation, VQA, and visual grounding.
@@ 3.1.1 Dataset Overall
The dataset covers six core SAR image analysis tasks: classification, fine-grained description, instance counting, spatial grounding, cross-modal identification, and referring. Based on the SARDet-100K dataset (Dai et al., 2024), it incorporates multimodal adaptations and enhanced language annotations from ten established SAR detection benchmarks such as AIR-SARShip(1.0&2.0)(Wang et al., 2019a), HRSID(Wei et al., 2020), MSAR(Chen et al., 2022), SADD(Zhang et al., 2021a), SAR-AIRcraf(Zhirui et al., 2023), ShipDataset(Wang et al., 2019b), SSDD(Zhang et al., 2021b), OGSOD(Wang et al., 2023), and SIVED(Lin et al., 2023). The dataset establishes image-text correspondences across six semantic categories (ships, tanks, bridges, ports, aircraft, and automobiles) through cross-modal representation learning, resulting in 2 million curated annotations.
@@ 3.1.2 Task Definition
(2) Fine-Grained Description: This task requires the VLM to identify SAR target categories and analyze their geometric attributes (morphological features and spatial orientations), evaluating the model’s ability to reason about SAR-specific spatial-geometric relationships.
@@
(4) Spatial Grounding: This task enables the VLM to understand and express spatial relationships (positions, distances, directions) among multiple targets in SAR images, which is crucial for military reconnaissance scenarios.
@@ 3.2 Task-Oriented Data Generation
Based on the characteristics of the six tasks, this study designs a multimodal dialogue data generation scheme, and its specific rules and implementation logic are as follows:
@@
Q&A Example:   User: "Please identify the target categories in the figure (candidate categories: {total category list})" AI: "The {specific category} targets can be identified in the figure"
@@
This task evaluates the model’s structured parsing of satellite imagery through comprehensive quality control. Following our dataset definitions, we filter images below 224×224 pixels and exclude targets with area ratio R < 1
@@
Q&A Example:   User: "Please describe the content of the satellite image in detail" AI: "In the figure, {size}{category} can be seen located at {orientation},"+"and there is also {size}{category} located at {orientation}"
@@
This task evaluates the model’s quantitative counting capabilities. We prefix 15 question templates with $[count]$ identifiers for task specification, while leveraging our unified bounding box format for structured outputs. The task supports extended expressions for multi-instance scenarios through coordinate serialization.
Q&A Example:   User: "[count] How many {category} are there in the image?"" AI: "There are 3 instances. {bbox}{bbox}{bbox}"
@@
This task assesses the model’s capability in expressing structural spatial relationships between multiple targets. Based on the established grid system, we quantify spatial relationships through relative distance metrics (proximal threshold: $(W_{img}+H_{img})/8$) and directional relationships (horizontal, vertical, diagonal orientations). The task comprises 15 spatial-relationship templates prefixed with $[grounding]$ identifiers, adhering to our unified spatial representation scheme.
Q&A Example:   User: "[grounding] Can you describe the spatial arrangement of various features in this SAR image?" AI: "In this SAR image, I can identify several objects and their locations: There is a {category-1} in the top right portion {bbox}. There are 2 {category-2} in the bottom left: one in the bottom left portion {bbox}, another in the middle left portion {bbox}. The {category-1} is far from the {category-2}."
@@
Q&A Example:   User: "[identify] What object is located in {bbox}?" AI: "This is a medium-sized {category} located in the top right portion of the image. It appears to be wide in shape."
@@
This task evaluates cross-modal correlation capabilities between natural language and image regions. Queries follow the pattern "Where is the {category}?", prefixed with $[refer]$ identifiers. The task outputs both precise bounding box coordinates and grid-based orientation descriptions, adhering to our unified spatial representation framework through nested parenthetical notation.
Q&A Example:   User: "[refer] Where is the {category}?" AI: "{bbox} (top right)."
@@ 3.3 Quantitative Analysis of Datasets
This study conducts quantitative analyses from two dimensions: category distribution and object morphology.
@@
The dataset exhibits three distinct morphological categories based on aspect ratio (AR): broad-bodied (AR $\leq$ 0.67), nearly square-shaped (0.67 $<$ AR $\leq$ 1.5), and tall-bodied (AR $>$ 1.5). For more distribution analysis, please refer to the Appendix .1.2.
@@ 4 Evaluation Method and Settings
This section aims to introduce the evaluation methods for six key tasks. These tasks encompass the model’s capabilities in multiple core areas such as information processing, target localization, and semantic understanding, and can reveal the model’s strengths and weaknesses from different dimensions.
@@ 5.2 Benchmark Evaluation
To verify the effectiveness and practicality of the SARChat-2M dataset, this study presents SARChat-Bench, a series of benchmark experiments on trained models which are based on this dataset. As shown in Table 3, sixteen mainstream visual-language models, such as Qwen2-VL(Wang et al., 2024), InternVL2.5(Chen et al., 2024), DeepSeekVL(Lu et al., 2024), Phi-3.5-vision(Abdin et al., 2024), GLM-Edge-V(GLM et al., 2024), mPLUG-Owl3(Ye et al., 2024), Yi-VL(Young et al., 2024) and LLaVA(Liu et al., 2023). In experiments of different tasks, each model exhibits diverse characteristics:
@@
(1) Instance Counting: This task requires VLMs to identify the number of specific objects in the image. Two leading model families achieve state-of-the-art performance: InternVL2.5 and Qwen-VL2, reaching accuracies of 74.14% and 72.79% respectively. However, the accuracy of most other models fall below 60%, highlighting both the challenging nature of the counting task and the dataset’s effectiveness in differentiating model capabilities.
@@
(2) Spatial Grounding: This task evaluates models’ capability in spatial localization. For abstract position descriptions, mPLUG-Owl3-7B achieves 100% accuracy, significantly outperforming other models. The mPLUG-Owl3 family maintains superior performance (>90%) in single-target localization, while other models achieve 80%-85%. However, in multi-target scenarios, most models’ accuracy drops to approximately 60%. These results suggest that accurate multi-target spatial information processing remains a crucial area for future model improvements.
@@
(3) Cross-Modal Identification: This task focuses on the model’s ability to build precise connections between visual information and other modal information. In this experiment, the process from image recognition to text description is mainly concerned. The experimental data shows that for both single-target and multi-target tasks, the accuracy rates of most models exceed 90%. Among them, the mPLUG-Owl3-7B model performs the best, with the accuracy rates of single-target and multi-target tasks reaching 99.27% and 99.51% respectively, fully demonstrating the powerful capabilities of large language models in cross-modal identification tasks.
@@
(4) Referring: This task challenges models to precisely locate objects in SAR images based on textual descriptions. Our experiments reveal significant performance gaps: models achieve less than 75% accuracy on single-target tasks and below 40% on multi-target scenarios. These results highlight the current limitations in cross-modal alignment, particularly in establishing precise text-to-object correspondences within SAR imagery.
@@
(5) Fine-Grained Description: This task requires the model to provide detailed feature and attribute descriptions of the objects in the image. The experiment shows that the model accuracy rates are in the range of 40%-63%. Among them, models with larger parameters such as Qwen2-VL-7B and InternVL2.5-8B perform outstandingly and can give more detailed and accurate descriptions. In contrast, other models with smaller parameter sizes perform poorly, indicating that the accuracy rate of the fine-grained description task is highly sensitive to the model’s parameter size.
@@
(6) Classification: The image classification task evaluates models’ ability to categorize images based on their content. According to the table data, regardless of parameter size, series such as InternVL2.5, mPLUG-Owl3, Qwen2-VL, and several other models achieve accuracy rates exceeding 96%. The performance of these VLMs demonstrates competitiveness with traditional vision classification models.

### v2

@@ Abstract
As a powerful all-weather Earth observation tool, synthetic aperture radar (SAR) remote sensing enables critical military reconnaissance, maritime surveillance, and infrastructure monitoring. Although Vision language models (VLMs) have made remarkable progress in natural language processing and image understanding, their applications remain limited in professional domains due to insufficient domain expertise. This paper innovatively proposes the first large-scale multimodal dialogue dataset for SAR images, named SARChat-2M, which contains approximately 2 million high-quality image-text pairs, encompasses diverse scenarios with detailed target annotations. This dataset not only supports several key tasks such as visual understanding and object detection tasks, but also has unique innovative aspects: this study develop a visual-language dataset and benchmark for the SAR domain, enabling and evaluating VLMs’ capabilities in SAR image interpretation, which provides a paradigmatic framework for constructing multimodal datasets across various remote sensing vertical domains. Through experiments on 16 mainstream VLMs, the effectiveness of the dataset has been fully verified. The project will be released at https://github.com/JimmyMa99/SARChat.
@@
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author. Xiayang Xiao and Haipeng Wang
@@ 1 Introduction
In recent years, deep neural networks, notably CNNs(LeCun et al., 1998) and ViTs(Dosovitskiy, 2020), have achieved remarkable progress in remote sensing data analysis, enhancing both processing efficiency and analytical accuracy. However, existing research mainly focuses on visual feature extraction, while lacking deep semantic parsing and reasoning capabilities (Li et al., 2024), limiting model applicability in complex scenarios.
With the advancement of Large-Language Models (LLMs), Vision-Language Models (VLMs), through integrating pre-training and instruction tuning, have demonstrated robust zero-shot learning and generalization in multimodal tasks Dai et al. (2023). This has inspired researchers to explore the deep integration of visual models with LLMs.
Although models designed for optical remote sensing images, like RSGPT (Hu et al., 2023) and GeoChat (Kuckreja et al., 2024), have shown preliminary achievements, they struggle to perform well in SAR applications. SAR images inherently pose significant interpretation challenges due to their scattering imaging mechanisms, characterized by blurred target edges, dispersed speckles, and orientation sensitivity. Meanwhile, existing SAR datasets primarily focus on visual recognition tasks (Kuckreja et al., 2024; Cheng et al., 2022; Zhang et al., 2023), leaving a critical shortage of large-scale, high-quality image-text alignment datasets. Both these intrinsic characteristics and data limitations impede the advancement of VLMs in the SAR domain.
@@
Current VLMs are primarily trained on conventional natural images without extensive fine-tuning for the SAR vertical domain. Despite their strong visual capabilities for natural images, these VLMs still have significant room for improvement in SAR image interpretation. Building upon the SARDet-100K dataset (Dai et al., 2024) with its rich SAR imagery and detection annotations, we construct SARChat-Bench-2M, a task-oriented SAR-specific image-text pair dataset, to address the insufficient SAR image interpretation capabilities of existing VLMs.
As shown in Figure 1, we present SARChat-2M, a large-scale multimodal conversational dataset for SAR images, and establish SARChat-Bench, a comprehensive multimodal task-oriented benchmark for the SAR domain. The SARChat-2M dataset contains approximately 2 million high-quality SAR image-text pairs across maritime, terrestrial, and urban scenarios, featuring fine-grained semantic descriptions and multi-scale resolutions (0.3-10 meters). The dataset supports major vision-language tasks such as image captioning, VQA(Visual Question Answering), visual localization, and object detection. To systematically evaluate model performance in these domains, we design six specific benchmark tasks in SARChat-2M: classification, fine-grained description, instance counting, spatial grounding, cross-modal identification, and referring. To validate the effectiveness of our dataset and benchmark, we conduct extensive experiments by fine-tuning 16 state-of-the-art VLMs of varying parameter scales, including InternVL2.5, DeepSeekVL, GLM-Edge-V, and the mPLUG-Owl family. Through training on SARChat-2M, these visual language models (VLMs) acquire comprehensive multi-task capabilities in SAR interpretation, as demonstrated by our systematic evaluation on SARChat-Bench.
The primary contributions of this paper are as follows:
  1. 1.
    The construction of SARChat-2M, the largest SAR remote sensing instruction-following dataset to date, comprising over 2 million high-quality image-text pairs across multi-scenario task-oriented dialogues, alleviating the knowledge scarcity of VLMs in the SAR domain.
  2. 2.
    The development of SARChat-Bench, a comprehensive SAR domain multimodal benchmark encompassing six core tasks (classification, description, counting, localization, recognition, and refering), enabling systematic evaluation of vision-language models through multi-dimensional assessment metrics.
  3. 3.
    It pioneers a research paradigm applicable to the SAR field, providing reference ideas for the construction of models in other remote-sensing vertical domains. The methods and processes adopted in data collection, annotation, as well as model training and evaluation in this study have good generality and extensibility.
@@ 2.1 VLMs for Remote Sensing
VLMs are capable of converting images into natural language descriptions and parsing the relationships between objects, demonstrating remarkable performance in tasks such as text-image retrieval, image captioning, and visual question answering. Recently, models like RemoteClip (Liu et al., 2024) have been applied to the field of remote sensing images, primarily focusing on cross-modal retrieval and zero-shot classification. However, these models have not addressed tasks such as image description generation and visual grounding. The RSGPT model has achieved text description and visual question answering for remote sensing images, but it has not expanded to tasks such as classification and detection. The GeoChat model has advanced multi-task conversational processing of high-resolution remote sensing imagery, including scene classification, visual question answering, multi-turn dialogue, visual grounding, and reference object detection. However, these models, including GeoChat, predominantly rely on optical remote sensing training data, leading to suboptimal performance in SAR-specific interpretation tasks. EarthGPT (Zhang et al., 2024) has extended the application of multimodal large language models to the remote sensing field through instruction tuning, but its performance in SAR image multi-task processing still needs improvement. Compared with natural images, the interpretation of SAR images is more challenging, which poses higher demands on the model’s processing capabilities and adaptability.
@@ 2.2 Remote Sensing Vision-Language Datasets
Remote sensing datasets are essential for intelligent interpretation models. Existing datasets such as UCM Captions Qu et al. (2016b), Sydney Captions Qu et al. (2016a), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), and RSVG (Zhan et al., 2023) provide preliminary resources for studying the correlation between remote sensing images and text. However, these datasets are limited not only in scale but also in modality, containing only optical images without SAR data, leaving SAR interpretation capabilities largely unexplored. Although large-scale datasets like MillionAID (Long et al., 2021), FMoW (Christie et al., 2018), and BigEarthNet (Sumbul et al., 2019) exist, they lack text-image pairs. The RS5M dataset (Zhang et al., 2023), containing 5 million image-text pairs, is still limited to optical images. The MMRS-1M dataset (Zhang et al., 2024), which covers optical, infrared, and SAR modes, has a very low proportion of SAR image-text data. Therefore, this paper constructs the SARChat-2M dataset, which focuses on SAR images and contains over 2 million image-text pairs, covering tasks such as classification, detection, caption generation, VQA, and visual grounding.
@@ 3.1.1 Dataset Overall
The dataset covers six core SAR image analysis tasks: classification, fine-grained description, instance counting, spatial grounding, cross-modal identification, and referring. Based on the SARDet-100K dataset (Dai et al., 2024), it incorporates multimodal adaptations and enhanced language annotations from ten established SAR detection benchmarks such as AIR-SARShip(1.0&2.0)(Wang et al., 2019a), HRSID(Wei et al., 2020), MSAR(Chen et al., 2022), SADD(Zhang et al., 2021a), SAR-AIRcraf(Zhirui et al., 2023), ShipDataset(Wang et al., 2019b), SSDD(Zhang et al., 2021b), OGSOD(Wang et al., 2023), and SIVED(Lin et al., 2023). The dataset establishes image-text correspondences across six semantic categories (ships, tanks, bridges, ports, aircraft, and automobiles) through cross-modal learning, resulting in 2 million curated annotations.
@@ 3.1.2 Task Definition
(2) Fine-Grained Description: The fine-grained description task focuses on both target category identification and geometric attribute analysis in SAR imagery. Beyond basic classification, it evaluates the VLM’s capability to extract detailed morphological features and spatial orientations, demonstrating the model’s proficiency in reasoning about SAR-specific spatial-geometric relationships.
@@
(4) Spatial Grounding: Spatial relationship interpretation enables comprehensive analysis of target distributions in SAR imagery, focusing on positions, distances, and directional relationships among multiple targets. This capability is particularly vital for military reconnaissance scenarios, where precise spatial understanding directly impacts operational effectiveness.
@@ 3.2 Task-Oriented Data Generation
Based on the characteristics of the six tasks, this study designs a multimodal dialogue data generation scheme. The specific rules and implementation logic are as follows, with detailed examples provided in the Appendix .1.4:
@@
Fine-Grained description evaluates the model’s structured parsing of satellite imagery through comprehensive quality control. Following our dataset definitions, we filter images below 224×224 pixels and exclude targets with area ratio R < 1% (Equation 1). Targets with aspect ratios exceeding 10:1 or out-of-bounds coordinates are removed. Size descriptions are categorized using area-ratio thresholds (small: <5%, large: >30%). We construct 40 interaction templates to accommodate multi-target scenarios.
@@
As a fundamental component of our visual reasoning system, this task focuses on evaluating the model’s quantitative counting capabilities. We designate 15 question templates with $[count]$ identifiers to specify the task requirements, while utilizing our unified bounding box format for structured output representation. The framework supports extended expressions for multi-instance scenarios through coordinate serialization.
@@
Spatial Grounding assessment evaluates the model’s proficiency in characterizing structural relationships among multiple target objects. Leveraging our established grid system, we quantify spatial relationships through two primary mechanisms: relative distance metrics (with proximal threshold defined as $(W_{img}+H_{img})/8$) and directional relationships (encompassing horizontal, vertical, and diagonal orientations). The framework incorporates 15 spatial-relationship templates, each prefixed with $[grounding]$ identifiers, conforming to our unified spatial representation scheme.
@@
Referring evaluates cross-modal correlation capabilities between natural language and image regions. Queries follow the pattern "Where is the {category}?", prefixed with $[refer]$ identifiers. The task outputs both precise bounding box coordinates and grid-based orientation descriptions, adhering to our unified spatial representation framework through nested parenthetical notation.
@@ 3.3 Quantitative Analysis of Datasets
The quantitative analysis in this study focuses on two key dimensions: category distribution and object morphological patterns.
@@
The dataset exhibits three distinct morphological categories based on aspect ratio (AR): broad-bodied (AR $\leq$ 0.67), nearly square-shaped (0.67 $<$ AR $\leq$ 1.5), and tall-bodied (AR $>$ 1.5). Detailed distribution analysis can be found in Appendix .1.2.
@@ 4 Evaluation Method and Settings
This section aims to introduce the evaluation methods for six tasks. These tasks evaluate the model’s core capabilities across information processing, target localization, and semantic understanding, providing multi-dimensional insights into visual-language model performance.
@@ 5.2 Benchmark Evaluation
To verify the effectiveness and practicality of the SARChat-2M dataset, this study presents SARChat-Bench, a series of benchmark experiments on trained models which are based on this dataset. As shown in Table 3, sixteen mainstream visual-language models, such as Qwen2-VL(Wang et al., 2024), InternVL2.5(Chen et al., 2024), DeepSeekVL(Lu et al., 2024), Phi-3.5-vision(Abdin et al., 2024), GLM-Edge-V(GLM et al., 2024), mPLUG-Owl3(Ye et al., 2024), Yi-VL(Young et al., 2024) and LLaVA-1.5(Liu et al., 2023). In experiments of different tasks, each model exhibits diverse characteristics:
@@
(1) Instance Counting requires VLMs to identify the number of specific objects in the image. Two leading model families achieve state-of-the-art performance: InternVL2.5 and Qwen-VL2, reaching accuracies of 74.14% and 72.79% respectively. However, the accuracy of most other models fall below 60%, highlighting both the challenging nature of the counting task and the dataset’s effectiveness in differentiating model capabilities.
@@
(2) Spatial Grounding evaluates models’ capability in spatial localization. For abstract position descriptions, mPLUG-Owl3-7B achieves 100% accuracy, significantly outperforming other models. The mPLUG-Owl3 family maintains superior performance (>90%) in single-target localization, while other models achieve 80%-85%. However, in multi-target scenarios, most models’ accuracy drops to approximately 60%. These results suggest that accurate multi-target spatial information processing remains a crucial area for future model improvements.
@@
(3) Cross-Modal Identification focuses on the model’s ability to build precise connections between visual information and other modal information. In this experiment, the process from image recognition to text description is mainly concerned. The experimental data shows that for both single-target and multi-target tasks, the accuracy rates of most models exceed 90%. Among them, the mPLUG-Owl3-7B model performs the best, with the accuracy rates of single-target and multi-target tasks reaching 99.27% and 99.51% respectively, fully demonstrating the powerful capabilities of large language models in cross-modal identification tasks.
@@
(4) Referring challenges models to precisely locate objects in SAR images based on textual descriptions. Our experiments reveal significant performance gaps: models achieve less than 75% accuracy on single-target tasks and below 40% on multi-target scenarios. These results highlight the current limitations in cross-modal alignment, particularly in establishing precise text-to-object correspondences within SAR imagery.
@@
(5) Fine-Grained Description requires the model to provide detailed feature and attribute descriptions of the objects in the image. The experiment shows that the model accuracy rates are in the range of 40%-63%. Among them, models with larger parameters such as Qwen2-VL-7B and InternVL2.5-8B perform outstandingly and can give more detailed and accurate descriptions. In contrast, other models with smaller parameter sizes perform poorly, indicating that the accuracy rate of the fine-grained description task is highly sensitive to the model’s parameter size.
@@
(6) Classification evaluates models’ ability to categorize images based on their content. According to the table data, regardless of parameter size, series such as InternVL2.5, mPLUG-Owl3, Qwen2-VL, and several other models achieve accuracy rates exceeding 96%. The performance of these VLMs demonstrates competitiveness with traditional vision classification models.